The Length and the breadth of a rectangular piece of land are 500 m and 300 m respectively. Find (i) Its area (ii) the cost of the land, if 1 m2 of the land costs ₹ 10,000.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

From the question it is given that,
Length of the rectangular piece of land = 500 m
Breadth of the rectangular piece of land = 300 m
Then,
(i) Area of rectangle = Length × Breadth
= 500 × 300
= 150000 m^2
(ii) Cost of the land for 1 m^2=₹10000
Cost of the land for 150000 m^2= 10000 × 150000
=₹1500000000

​Thanks